Question 43) L12:) 330003-21 + jou X [-2 -1 ][x₂] li Derive the transfer function y(s) = Gis)u(s) 4) Using MATLAB, determine the step
Answer #1

2:5] Given *=[8 0611-60] yo [0 ] [%] + [0] 4 & [40) = Growcss Transfer function - Go) = C(51-45/B +D (51 -A) = (* :]-13 -] 6from 0.2) Go)= Co Je [ ][1]** 2x1 GCS) - k Transfer function 5+5+2 Step Response of yos)- GS) ucs) :- LIS) - (son (695 - (06By Iovesse Laplace Transformer yet) a ovocat sincet) Astep Response of system7 Editor - Untitled.m Command Window Untitled. m x + 1 - sys = tf([1 0],[1 1 2]); Figure 1 File Edit View Insert Tools Deskto4) 0) At [ X(t) = e X (0) + Alt-) Bult) do Since xco)=0 X(t) = Alt- ) . Bult de Ho | xce) - ot-re acc c ó - $ y =, (4) where2015+ [cost] x + usinąt] to zostsinnt At from 6 » X(t)= lealt-2) B(C) ulos de uit) a step with Amplitude 6 - ues) = after Sim